STAP 5.) A pulley is rotating at an angular velocity of 32 revolutions per minute. The radius of the pulley is 0.3 m. a.) How what is the angular distance the pulley rotates in 5 sec? b.) What is the linear distance the pulley rotates in 5 sec? c.) If a motor takes 30 sec to speed up the pulley to 82 revolutions per minute, what was the angular acceleration the motor applied? d.) How far will the pulley have traveled while accelerating?STAP Rotation Homework 3 1.) Calculate the force of friction that keeps an 80 kg person sitting on the edge of a yground merry-go-round when the person sits 2 m from the center, and has a tangential speed of 3 m/s 2.) To ti ghten a bolt, you push with a force of 80 N perpendicular to the handle of a at a distance of 0.25 m from the center of the bolt. a.) What torque are you applying to the bolt? wrench b.) You move your hand inward so that you are only 0.1 m from the end of the bolt. Show that to achieve the same torque, you would need a force of 200N. 3.) Calculate the torque produced by a 50 N force applied at an angle of 75° at the end of a 1m long wrench. 4.) A hockey puck of mass m revolves on an icy surface in a circle at speed v at the end of a horizontal string of length L. The tension in the string is T a.) Write the equation for the centripetal force, and substitute the values of T and L appropriately b.) Show that the mass of the puck is 5 kg when the length of the string (L) is 2 m, the tension (T) is 10 N, and the tangential speed (v) is 2 m/s
Answer #1

1)

Force of friction, Fs = m*v^2/r

= 80*3^2/2

= 360 N

2)

a)

a) Torque = r1*F1*sin(theta)

= 0.25*80*sin(90)

= 20 N.m

b) Torque = r2*F2*sin(90)

F2 = Torque/r2

= 20/0.1

= 200 N

3) Torque = r*F*sin(theta)

= 1*50*sin(75)

= 48.3 N.m

4)

a) cenetripetal force = m*v^2/L

T = m*v^2/L

b) T = 5*2^2/2

= 10 N

5)

a) theta = w*t

= 32*(2*pi/60)*5

= 16.8 radaians

b) d = r*theta

= 0.3*16.8

= 5.04 m

c) alfa = (wf - wi)/t

= (82 - 32)*(2*pi/60)/30

= 0.174 rad/s^2

d) theta = wi*t + (1/2)*alfa*t^2

= 32*(2*pi/60)*30 + (1/2)*0.174*30^2

= 179 radians